Fermi National Laboratory
 
Operation Summary for March 9, 2012 - March 12, 2012
 

NuMI reported receiving 7.84E18 protons on target for the period from 3/5/12 to 3/12/12.

Friday March 9

The day shift began with the Booster off for an access to investigate a sump pump high-level alarm, with Operations monitoring the Linac and Main Injector (MI), and with the NuMI and MiniBooNE experiments off do to the Booster access.

During the day shift, at 8 AM, Neutron Therapy Facility (NTF) personnel prepared to treat one patient later in the morning. Also at 8 AM, a Booster RF technician took a look at an RF station (BRF2). The Booster came out of access, reported it was a stuck float, and resumed sending beam by 8:34 AM. Operations began sending beam to the Neutrino Area SeaQuest experiment E-906 around 8:45 AM. A Booster person began an aperture scan study at 8:50 AM. At 9:34 AM, the MiniBooNE beam permit tripped off due to bad beam positions; it reset. Operations established beam to the Meson FTBF experiment T-1008 at 10:21 AM. At 1:01 PM, Operation halted the Linac beam to allow techs to work on a RF station (LRF3). Operators loaded an energy conservation time line into the Time Line Generator (TLG) at 1:08 PM. Linac techs completed their work on LRF3 and Linac resumed sending beam at 1:38 PM. At 2:48 PM, a NuMI kicker (I:KPS6N) tripped off; a Kicker person investigated and reported that a cable needed to be reterminated. Operators loaded an energy conservation time line at 3:25 PM. Operators prepared the Neutrino Area enclosures NM3 and NM4 for an access at 3:29 PM. Work on I:KPS6N ended and MI resumed sending beam at 3:55 PM.

During the evening shift, a NuMI person reported at 4:48 PM that while NuMI was down due to a their kicker problem the MI-62 LCW system lost about 100 gallons of water, but when they resumed taking beam, the leak went away; Operators created an alarm to track this problem. NuMI resumed taking beam at 5:06 PM. The Booster studies ended at 5:19 PM. At 6:45 PM, Operators reported that they had completed an MI tune, injection closure, and made phase and energy changes. Operations reported at 10:16 PM that they were still struggling with MI beam position problems.

Saturday March 10

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-1008.

During the midnight shift, Operations disabled beam to T-1008 at 1:09 AM. Operators loaded protons into the Recycler at 3:40 AM. Operations established beam to T-1008 at 7:41 AM.

During the day shift, at 9:01 AM, the MiniBooNE beam permit tripped off for the first of five times in six hours due to bad beam positions; it reset.

During the evening shift, at 9:23 PM, LRF3 tripped off and would not reset; Operations called in Linac personnel. Operators loaded an energy conservation time line at 9:39 PM. Linac resumed sending beam at 10:12 PM.

Sunday March 11

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-1008.

During the midnight shift, at 6:34 AM, LRF4 tripped off and would not reset; Operations called in Linac personnel.

During the day shift, Linac personnel replaced a fuse for LRF4 and Linac resumed sending beam around 9:15 AM. Operations reported at 9:40 AM that while the Linac beam was off, NuMI cooled down and its LCW leak returned; the leak stopped when NuMI resumed taking beam.

During the evening shift, at 4:43 PM, Operations reported problems with Booster RF sparking; investigations revealed that BRF19 was the cause; Operators made adjustments. Operations completed a Linac and Booster tune 8:09 PM. Operators loaded protons into Recycler at 9:21 PM. Operations received a call at 10:51 PM from the Cryo System Coordinator about the MTA cryo system; Operators investigated and found the +5 volts missing from an IO crate. The Cryo Coordinator called in a tech who had the problem repaired by 11:45 PM.

Monday March 12

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-1008.

During the midnight shift, Operations completed a 750 KeV line tune by at 12:06 AM. Operations completed a Booster injection tune by 12:52 AM. At 6:05 AM, the H- Source tripped off; it reset.

Operations Coordinator (OpsCo) Plans for Today

The OpsCo agrees with the plans of MI personnel to conduct a g-2 (g minus 2) study. NuMI personnel would like to access MI-62 to search for a possible LCW leak when an opportunity arises. SeaQuest would also like to make an access into NM4 when an opportunity arises.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/12/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory